OpenStudy (anonymous):

y-(-10-c)=z solve for c

OpenStudy (anonymous):

y-(-10-c)=z Distribute the -1: y + 10 + c = z Subtract y+10 from both sides: c = z - y - 10

OpenStudy (anonymous):

thanks